OpenStudy (maddy1251):

Need help?? I have the problem -3tan^2(2x)+1=0 And I am having trouble solving it.. This is where I've gotten so far. -3tan^2 (2x)=1 -3tan^2(x)=1/2 This is where I get stuck lol can anyone help?

OpenStudy (maddy1251):

@robtobey

OpenStudy (dumbcow):

First off, you should have -1 on right side after step 1 Second, the "2" in tan(2x) is attached to the tan function and cannot be divided on both sides

OpenStudy (maddy1251):

@dumbcow You're right.. \[-3\tan^{2}2x=-1\]

OpenStudy (maddy1251):

Then what?

OpenStudy (dumbcow):

Originally think of the equation as \[-3u^2 +1 = 0\] where i replaced tan function with variable u it can be easier to solve if you are not confused by the trig

OpenStudy (maddy1251):

So -3u^2=-1 Would you then factor -3u^2?

OpenStudy (dumbcow):

how would you factor that? remember its only 1 term

OpenStudy (maddy1251):

u^2= -1/-3 u^2=1/2 sqrt(^2)=sqrt(1/2)?

OpenStudy (maddy1251):

Tht should be a 3,, not 2.

OpenStudy (dumbcow):

very good

OpenStudy (maddy1251):

My bad. Okay, so would it plus and minus?

OpenStudy (dumbcow):

Not lets put back in the tan .... yes plus minus \[u = \pm \frac{\sqrt{3}}{3}\] \[\tan(2x) = \frac{\sqrt{3}}{3}\] You must use inverse tangent to cancel the tan function to get to the "x" \[2x = \tan^{-1} (\frac{\sqrt{3}}{3})\]

OpenStudy (dumbcow):

use calculator or unit circle to find angles sorry i for got plus minus ... you also look up inverse tan of neg sqrt3/3

OpenStudy (dumbcow):

have you found the angles?

OpenStudy (dumbcow):

ok this is what you should get from your calculator (using degrees) \[\tan^{-1} \pm \frac{\sqrt{3}}{3} = \pm 30\] Note the period for tangent is 180 so if you need all solutions, add multiples of 180 \[2x = \pm 30 + 180n\] Finally solve for x by dividing by 2 \[x = \pm 15 +90n\]

OpenStudy (maddy1251):

Thanks @dumbcow ! I actually got x= +/- 15 but forgot the 90n. Thank you! I don't feel so dumb now thanks to you.

OpenStudy (maddy1251):

I had forgot bout the period of the tangent being 180

OpenStudy (dumbcow):

haha no problem , glad to help